very unsure how to solve this or what numbers to plug into the equations there are so many combinations. If I had to guess I would guess "No Solutions"
Solve the system Write your answer using parentheses and no spaces. Example (2,0,-3) or write "No Solution" for questions with no solution.
x + y + z = 6
2x - y + 2z = 6
-x + y + 3z = 10

Answer by TutorDelphia(193)

Solve using the elimination method rather than guessing: First elimination one variable, then another
x + y + z = 6
plus
-x + y + 3z = 10
2y+4z=16

2x - y + 2z = 6
plus 2 times
-x + y + 3z = 10
-2x+2y+6z=20
which gives
y+5z=26
We already had:
2y+4z=16
multiply y+5z=26 by -2
-2y-10z=-52
-6z=-36
z=6
Now that we know z=6 sub it in 2y+4z=16
2y+24=16
2y=-8
y=-4
x + y + z = 6
x+-4+6=6
x=4
(4,-4,6)